Chapter 5 P33 of Survey for Accounting 5th edition I am completely lost on filling out the chart from the question.

Page 202

The Brick Company had cash sales of $280,000 for 2018, its first year of operation. On April 2, the company purchased 210 units of inventory at $390 per unit. On September, an additional 160 units were purchased for $425 per unit. The company had 110 units on hand at the end of the year. The company's income tax rate is 40%. All transactions are cash transactions.

  1. The preceding paragraph describes five accounting events: (1) a sales transaction, (2) the first purchase of inventory, (3) a second purchase of inventory, (4) the recognition of cost of goods sold expense, and (5) the payment of income tax expense. Record the amounts of each event in horizontal statements models like the following ones, assuming first a FIFO and then a LIFO cost flow.

Effect of Events on Financial Statements

Panel 1: FIFO Cost Flow

Event No.

Balance Sheet

Income Statement

Statement of Cash Flows

Cash + Inventory = Ret. Earn.

Rev. – Exp. = Net Income

1

2

3

4

5
Answer #1

Under FIFO Method, Ending inventory are from last lot 110 units @ 425 = 46750

COGS = All purchases - Ending inventory

= 210*390 + 160*425 - 46750

= 81900 + 68000 - 46750

= 103150

Profit = Revenue - COGS

= 280000 - 103150

= 176850

Income tax = 176850*30%

= 53055

Net income = 176850 - 53055

= 123795

Event No. Balance Sheet Income Statement Statement of Cash Flows
Cash + Inventory = Ret. Earn. Rev. – Exp. = Net Income
1 Cash will be increased by 280000 Revenue will be increased by 280000 cash inflow from operating activities will be increased by 280000
2 inventory will be increased by 81900 & cash will be decreased 81900 NIL cash will be decreased by 81900
3 inventory will be increased by 68000 & cash will be decreased 68000 NIL cash will be decreased by 68000
4 NIL COGS will be recorded as expense, which decreased revenue 103150 NIL
5 Cash will be decreased by 53055 & from liability side it reduces its current liability it reduced its net income by 53055 cash will be decreased by 53055

Under LIFO Method, Ending inventory are from last lot 110 units @ 390 = 42900

COGS = All purchases - Ending inventory

= 210*390 + 160*425 - 42900

= 81900 + 68000 - 42900

= 107000

Profit = Revenue - COGS

= 280000 - 107000

= 173000

Income tax = 173000*30%

= 51900

Net income = 173000 - 51900

= 121100

Event No. Balance Sheet Income Statement Statement of Cash Flows
Cash + Inventory = Ret. Earn. Rev. – Exp. = Net Income
1 Cash will be increased by 280000 Revenue will be increased by 280000 cash inflow from operating activities will be increased by 280000
2 inventory will be increased by 81900 & cash will be decreased 81900 NIL cash will be decreased by 81900
3 inventory will be increased by 68000 & cash will be decreased 68000 NIL cash will be decreased by 68000
4 NIL COGS will be recorded as expense, which decreased revenue 107000 NIL
5 Cash will be decreased by 51900 & from liability side it reduces its current liability it reduced its net income by 51900 cash will be decreased by 51900

Please check with your answer and let me know.